OpenStudy (anonymous):

√20=_____√5 How do I solve this && what would the answer be im confused help me plzzz

OpenStudy (anonymous):

You're simplifying radicals

OpenStudy (anonymous):

A radical can split up by factors and you can also radicals of the same n-th root together

OpenStudy (anonymous):

\[\sqrt{2}\times \sqrt{3}=\sqrt{6}\]

OpenStudy (anonymous):

??? huh

OpenStudy (anonymous):

\[\sqrt{80}= \sqrt{16}\times \sqrt{5}\]

OpenStudy (anonymous):

All they're asking is to simplify this radical

OpenStudy (anonymous):

And a radical can be split up by its factors, the example of the square root of 80 16 and 5 are factors of 80 because 16x5=80

OpenStudy (anonymous):

So thats the answr

OpenStudy (anonymous):

so we can just write it as the square root of 16 times the square root of 5 but the square root of 16 is just 4 so we put it on the outside and write like \[4\sqrt{5}\]

OpenStudy (anonymous):

so thats the answr

OpenStudy (anonymous):

and they started with 20 and they split it up with the \[\sqrt{20}= \sqrt{5} \times what number?\]

OpenStudy (anonymous):

??? im confused

OpenStudy (anonymous):

ok do you understand that the \[\sqrt{20}\] can be split up into its factors?

OpenStudy (anonymous):

so another way of writing \[\sqrt{20}\] would be \[\sqrt{10}\times \sqrt{2}\]

OpenStudy (anonymous):

so thats the answe.. i need hlped but i want the answer cuz i have atime limite to be done

OpenStudy (anonymous):

well \[\sqrt{20}=\sqrt{4}\times \sqrt{5} \]

OpenStudy (anonymous):

\[\sqrt{4}\] is the answer but thats just 2 <---answer, but you should invest some time in understanding this

OpenStudy (anonymous):

try this one plzzz √80=_____√5

OpenStudy (anonymous):

I did that up there^^^^^^^